Assume that a plant operates 2000 hours per year and the demand rate for parts is 100,000 units per year. The circulation time for each kanban container is 24 hours. a) How many kanban containers are needed for a container size of 100 parts? b) What would be the effect of reducing the container size to 60 parts? c) What is the takt time for this process? d) What takt time is needed for 80,000 units per year?
Answer #1

Given,

Working time(w) = 2000 hours per year

Demand(d) = 100,000 units per year

Cycle time(t) = 24 hours= 24/2000 = 0.012 year

Now,

a) container size(c)=100

No. of containers needed = d*t/c = 100,000*0.012/100 = 12 containers

b) container size(c)=60

No. of containers needed = d*t/c = 100,000*0.012/60 = 20 containers

c) Takt time = working hours/demand= 2000/100,000= 0.02 hours per unit

d) demand=80000 units

Takt time = 2000/80,000= 0.025 hours per unit​​​​​​